1) Calculate the noise power and noise voltage generated by a 1M ohm resistor at 37 C over a 1Mhz bandwith?

2) Three messages A,B, C have probilities 0.4, 0.3 and 0.3 respectively of their occurence. If total 100 messages are sent out calculate the total information carried by them. Also calculate entropy of the signal?


1
Expert's answer
2021-02-15T17:39:59-0500

1)


"P=kTB\\P=(1.38\cdot10^{-23})(273+37)(10^6)\\P=4.28\cdot10^{-15}\ W"

"V=\sqrt{4kTBR}\\V=\sqrt{4(1.38\cdot10^{-23})(273+37)(10^6)(10^6)}\\V=1.31\cdot10^{-4}\ V"

2) The entropy of the signal


"H=0.4\log_2{\frac{1}{0.4}}+0.3\log_2{\frac{1}{0.3}}+0.3\log_2{\frac{1}{0.3}}\\H=1.57\frac{bits}{message}"

The total information carried by them


"R=1.57(100)=157\ bits"